OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. Choose one answer. a. y = x + 2 b. y = -x - 2 c. y = -x + 4 d. y = x - 4

OpenStudy (anonymous):

@AbdullahM

OpenStudy (dan815):

the slope must be the same

OpenStudy (dan815):

-1 is the slope now write the equation of a line with slope -1 and that goes through the point -3,1

OpenStudy (abdullahm):

parallel lines have the same slope, while perpendicular lines have the "oppsite reciprocal" slope which is another way to say when multiplied they equal to -1

OpenStudy (abdullahm):

So, which answer choices has -1 as the slope?

OpenStudy (anonymous):

b and c

OpenStudy (dan815):

which one does your point satisfy?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

would it be b?

OpenStudy (dan815):

why do u say that

OpenStudy (anonymous):

Im not sure why. But i think i's right.

OpenStudy (anonymous):

its

OpenStudy (dan815):

xD

OpenStudy (anonymous):

so how would i solve it?

OpenStudy (anonymous):

i know that its b or c

OpenStudy (abdullahm):

Good job on eliminating A and D

OpenStudy (abdullahm):

The way to find out between B and C is to plug in the point (-3,1)

OpenStudy (anonymous):

show me your work so i can understand it a bit better?

OpenStudy (anonymous):

i can do another one of these problems after to make sure i understand it

OpenStudy (abdullahm):

For example: b. y = -x - 2 b. 1=-(-3) -2 (-3,1) is in the form of (x,y) so we replace the x and y with those values

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

1= 3-2

OpenStudy (abdullahm):

so that means B is our answer since 1=1